### v2

# Simultaneous false discovery bounds for invariant causal prediction

Abstract
Invariant causal prediction (ICP, Peters et al., (2016)) provides a novel way to identify causal predictors of a response by utilizing heterogeneous data from different environments. One advantage of ICP is that it guarantees to make no false causal discoveries with high probability. Such a guarantee, however, can be too conservative in some applications, resulting in few or no discoveries. To address this, we propose simultaneous false discovery bounds for ICP, which provides users with extra flexibility in exploring causal predictors and can extract more informative results. These additional inferences come for free, in the sense that they do not require additional assumptions, and the same information obtained by the original ICP is retained. We demonstrate the practical usage of our method through simulations and a real dataset.
1 Introduction
Discovering causal predictors of the response of interest is usually the primary goal of scientific research. Based solely on observational data, the causal relationship might be unidentifiable. In such cases, heterogeneous data from different environments can be helpful for further identifiability and thus is a valuable source for causal discovery. Under the multi-environments setting, Peters et al., (2016) proposed a novel method, called invariant causal prediction (ICP), to identify causal predictors of a response by exploiting the invariance of the conditional distribution of the response given its causal predictors across environments. Compared to other causal discovery algorithms, one main advantage of ICP is that it provides a statistical confidence guarantee for its output $\widehat{S}^{\text{ICP}}$:
$P\left(\widehat{S}^{\text{ICP}}\subseteq S^{*}\right)\geq 1-\alpha,$
where $\alpha\in(0,1)$ is a nominal level and $S^{*}$ denotes the set of true causal predictors. That is, ICP ensures that all its discoveries are true causal predictors with a probability larger than $1-\alpha$. This is also known as the familywise error rate control guarantee.
The original ICP approach has been generalized to non-linear setting (Heinze-Deml et al.,, 2018), sequential data (Pfister et al.,, 2019), and transformation models (Kook et al.,, 2023). All these methods inherit the familywise error rate control guarantee. However, controlling the familywise error rate can be too conservative in many applications, especially when causal predictors are highly correlated with non-causal ones. In such cases, ICP may result in few or no causal discoveries. To address this, Heinze-Deml et al., (2018) proposed the so-called defining sets, which are the smallest sets guaranteeing to contain at least one true causal predictor. They are especially useful in the case where ICP returns an empty set.
In this paper, we propose a method to obtain simultaneous false discovery bound (Genovese and Wasserman,, 2006; Goeman and Solari,, 2011; Goeman et al.,, 2021) for ICP. Specifically, let $\mathcal{N}^{*}\subseteq[m]=\{1,\dots,m\}$ be the index set of non-causal predictors among $m$ predictors, the simultaneous false discovery bound is a function $t_{\alpha}:2^{[m]}\rightarrow\{0,\dots,m\}$, where $2^{m}$ denotes the power set of $[m]$, such that
$P(|R\cap\mathcal{N}^{*}|\leq t_{\alpha}(R)\text{ for any }R\subseteq[m])\geq 1-\alpha.$
As a result, one can freely check any set $R$ of interest, and a high probability upper bound of the false discoveries can be obtained by $t_{\alpha}(R)$. This contains the familywise error rate control and the defining sets of Heinze-Deml et al., (2018) as special cases. In particular, by searching for the largest set $R$ whose corresponding $t_{\alpha}(R)$ is 00, we obtain a set with familywise error rate control; by searching for the smallest sets whose corresponding $t_{\alpha}(R)\leq|R|-1$, we obtain all defining sets.
Extra causal information can be extracted using the simultaneous false discovery bound. As a quick illustration, for a simulated dataset with nine predictors shown in Section 4.1, no information is obtained by the original ICP approach as it returns zero discoveries. However, more causal information can be extracted by using our simultaneous false discovery bounds. For example, we can know that with a probability larger than $0.95$, there is at least one causal predictor in $\{X_{3},X_{7}\}$, at least two causal predictors in $\{X_{3},X_{5},X_{7},X_{8}\}$, and at least four causal predictors in all nine predictors. More simulations and a real data application can be found in Section 4.1.
To obtain the simultaneous false discovery bound for ICP, we proceed by considering a multiple testing problem (see (6)) that directly tests causal predictors. We show that the original ICP approach is equivalent to directly comparing certain p-values (see (8)) to the significance level. Then, by generalizing the null hypothesis (6) and p-value (7) to (10) and (11), respectively, we propose a simultaneous false discovery upper bound (12). This upper bound can be seen as a special case of the closed testing method by Goeman and Solari, (2011), but in our case, no closed testing adjustment is needed due to the specific form of the p-values. In comparison to the original ICP approach, our method provides additional causal information at the cost of testing more hypotheses, without introducing any additional statistical assumptions. We also discuss the idea of controlling the false discovery rate for ICP.
2 A brief recap of invariant causal prediction
For the sake of simplicity, we use linear models (Peters et al.,, 2016) for illustration. Consider the following linear structural equation model:
$Y^{e}\leftarrow\beta_{1}X^{e}_{1}+\cdots\beta_{m}X^{e}_{m}+\epsilon^{e}=\beta_{S^{*}}^{T}X^{e}_{S^{*}}+\epsilon^{e},$ (1)
[…]
The key idea of ICP is to test whether the conditional distribution of $Y^{e}$ is invariant across environments given predictors $X^{e}_{S}$ for some $S\subseteq[m]$. This leads to the following null hypothesis:
$H_{0,S}(\mathcal{E}):$ there exists a distribution Fϵ and a vector $\text{there exists a distribution $F_{\epsilon}$ and a vector }\gamma\in\mathbb{R}^{m}\text{ with support }S\text{ such that }$ (2)
[…]
Under the assumptions that there is no latent variable and $\mathcal{E}$ does not contain the environment where $Y$ is intervened, it is clear that $H_{0,S^{*}}(\mathcal{E})$ is true. However, there may be other sets than $S^{*}$ for which $H_{0,S}(\mathcal{E})$ is true, causing an identifiability issue. To this end, Peters et al., (2016) defined the so-called identifiable causal predictors under $\mathcal{E}$ as follows:
$S(\mathcal{E})=\bigcap_{S:H_{0,S}(\mathcal{E})\text{ is true }}S.$ (3)
Note that $S(\mathcal{E})\subseteq S^{*}$ as $H_{0,S^{*}}(\mathcal{E})$ is true.
ICP estimates the causal predictors by using the sample version of $S(\mathcal{E})$ in two steps: (i) For every $S\subseteq[m]$, test $H_{0,S}(\mathcal{E})$ at level $\alpha\in(0,1)$. (ii) Obtain the selected causal predictors by
$\widehat{S}^{\text{ICP}}(\mathcal{E})=\underset{S:H_{0,S}(\mathcal{E})\text{ is not rejected }}{\bigcap}S.$ (4)
The familywise error rate guarantee of $\widehat{S}^{\text{ICP}}(\mathcal{E})$ holds at level $\alpha$ because
$P\left(\widehat{S}^{\text{ICP}}(\mathcal{E})\subseteq S^{*}\right)=P\left(\underset{S:H_{0,S}(\mathcal{E})\text{ is not rejected }}{\bigcap}S\subseteq S^{*}\right)\geq P(H_{0,S^{*}}(\mathcal{E})\text{ is not rejected })\geq 1-\alpha.$ (5)
3 Simultaneous false discovery bounds for ICP
3.1 A multiple testing formulation of ICP
Instead of directly estimating $S(\mathcal{E})$ as in (4), another natural way is to form a multiple testing problem:
$H^{*}_{0,i}:i\in\mathcal{N}^{*},\quad i\in[m],$ (6)
where $\mathcal{N}^{*}=[m]\setminus S^{*}$. To this end, we propose the following p-value for $H^{*}_{0,i}$:
$p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S},$ (7)
where $p_{S}$ is a valid p-value for testing $H_{0,S}(\mathcal{E})$ (see (2)), that is, $P_{H_{0,S}(\mathcal{E})}(p_{S}\leq c)\leq c$ for any $c\in(0,1)$. Note that it can be obtained in the first step of ICP. All p-values in this paper depend on the environment set $\mathcal{E}$, but we omit this dependence for simplicity of notation. The validity of $p^{*}_{i}$ is shown in the following proposition.
Proposition 3.1.
For any $c\in(0,1)$, $P_{H^{*}_{0,i}}(p^{*}_{i}\leq c)\leq c$.
To connect to the original ICP approach, we need to introduce a slightly more complicated p-value. For a given $\tau\in(0,1)$, let
$\tilde{p}^{*}_{i}(\tau)=\max\{p^{*}_{i},\Pi_{S\subseteq[m]}\mathbbm{1}_{p_{S}\leq\tau}\}.$ (8)
Note that $\tilde{p}^{*}_{i}(\tau)$ is also valid because $\tilde{p}^{*}_{i}(\tau)\geq p^{*}_{i}$. Then, as the following proposition shows, the discovery set obtained by using ICP is equivalent to the one obtained by directly comparing $\tilde{p}^{*}_{i}(\alpha)$ to the significance level $\alpha$.
Proposition 3.2.
[…]
Hence, $\tilde{S}(\mathcal{E})$ possesses the familywise error rate control guarantee (see (5)). In fact, the discovery set
$\widehat{S}(\mathcal{E})=\{i:p^{*}_{i}\leq\alpha\}$ (9)
obtained by directly comparing $p^{*}_{i}$ to the significance level also controls the familywise error rate, as shown in the following proposition.
Proposition 3.3.
For $\alpha\in(0,1)$, we have $P\left(\widehat{S}(\mathcal{E})\subseteq S^{*}\right)\geq 1-\alpha.$
Therefore, both $\widehat{S}(\mathcal{E})$ and $\widehat{S}^{\text{ICP}}(\mathcal{E})$ (equivalent to $\tilde{S}(\mathcal{E})$) control the familywise error rate, and using $\widehat{S}(\mathcal{E})$ seems better because $p^{*}_{i}\leq\tilde{p}^{*}_{i}(\alpha)$. However, the scenario where $p^{*}_{i}<\tilde{p}^{*}_{i}(\alpha)$ only occurs when $p_{S}\leq\alpha$ for all $S\subseteq[m]$. In such cases, we have $\widehat{S}^{\text{ICP}}(\mathcal{E})=\emptyset$ and $\widehat{S}(\mathcal{E})=[m]$, which appears not meaningful. In practice, $\widehat{S}(\mathcal{E})$ and $\widehat{S}^{\text{ICP}}(\mathcal{E})$ are generally equivalent.
3.2 Simultaneous false discovery bounds for ICP
Based on the multiple testing formulation (6), we propose simultaneous false bounds by considering all intersection hypotheses. Specifically, for a given $S\subseteq[m]$, by generalizing (6) and (7), we consider null hypothesis
$H^{*}_{0,S}:i\in\mathcal{N}^{*}\text{ for all }i\in S$ (10)
and p-value
$p^{*}_{S}=\max_{I\subseteq[m]\setminus S}p_{I}.$ (11)
By using a similar argument as in Proposition 3.1, one can see that $p^{*}_{S}$ is valid.
Proposition 3.4.
For any $c\in(0,1)$, $P_{H^{*}_{0,S}}(p^{*}_{S}\leq c)\leq c$.
For any set $R\subseteq[m]$, let $t_{\alpha}(R)$ be the size of the largest subset of $R$ whose corresponding p-value is larger than $\alpha$. That is,
$t_{\alpha}(R)=\max\{|I|:I\subseteq R\text{ and }p^{*}_{I}>\alpha\}.$ (12)
Then, $t_{\alpha}(R)$ is a simultaneous false discovery upper bound, as shown in the following Theorem.
Theorem 3.1.
Let $\alpha\in(0,1)$, then
$P(|R\cap\mathcal{N}^{*}|\leq t_{\alpha}(R)\text{ for any }R\subseteq[m])\geq 1-\alpha.$ (13)
Based on the simultaneous guarantee (13), one can freely check any set $R$ of interest, and a high probability upper bound (12) on the false discoveries can be calculated. Equivalently, $|R|-t_{\alpha}(R)$ is a simultaneous true discovery lower bound:
$P(|R\cap S^{*}|\geq|R|-t_{\alpha}(R)\text{ for any }R\subseteq[m])\geq 1-\alpha.$ (14)
The simultaneous upper bound (12) is the same as the upper bound proposed by Goeman and Solari, (2011) for closed testing (Marcus et al.,, 1976). In fact, our method can be seen as a special case of closed testing. Compared to the standard closed testing procedure, the difference is that no closed testing adjustments are needed in our case. This is due to the specific form of $p^{*}_{S}$. In particular, if $p^{*}_{S}\leq\alpha$, we must have $p^{*}_{S^{\prime}}\leq\alpha$ for all $S\subseteq S^{\prime}$. Thus, if a hypothesis is locally rejected, it must be rejected by closed testing, so there is no need to implement closed testing adjustments.
At last, we mention that calculating the false discovery bound (12) can be computationally intractable for large $|R|$ (or large $m$). In fact, computation is also a fundamental issue for the original ICP approach, which requires testing $2^{|m|}$ hypotheses. Some tricks are suggested to deal with this issue, including first reducing the size of the potential causal predictors by using a pre-screening method such as Lasso. See Peters et al., (2016) for more discussions.
3.3 A discussion about false discovery rate control for ICP
With a multiple testing formulation (6) and valid p-values (7), an alternative strategy to obtain a less conservative result than $\widehat{S}^{\text{ICP}}(\mathcal{E})$ is to control the commonly used false discovery rate (Benjamini and Hochberg,, 1995; Benjamini and Yekutieli,, 2001; Lehmann and Romano,, 2022). Due to Proposition 3.3, however, directly comparing $p^{*}_{1},\dots,p^{*}_{m}$ to the significance level already controls the familywise error rate, so there is no point in further applying false discovery rate control procedures on these p-values. This observation is interesting because, typically, multiple testing corrections (e.g., Bonferroni correction) on raw p-values are required to control the familywise error rate. Thus, it is natural to ask: Can we obtain smaller valid p-values than $p^{*}_{1},\dots,p^{*}_{m}$? If so, one may apply false discovery rate control procedures to obtain more discoveries than ICP. However, in the following, we argue that this approach does not seem promising.
Specifically, we show that if some $p^{\prime}_{i}$ satisfying
${\mathbb{P}}(p^{\prime}_{i}\leq p^{*}_{i})=1\quad\text{and}\quad{\mathbb{P}}_{H^{*}_{0,i}}(p^{\prime}_{i}<p^{*}_{i})>0,$ (15)
then it is not valid in general. In particular, suppose that we have an ideal test for $H_{0,S}(\mathcal{E})$ that yields p-values $p_{S}\sim U(0,1)$ for true nulls and $p_{S}=0$ for false nulls. Consider a setting where the only true null hypothesis is $H_{0,S^{*}}(\mathcal{E})$. This is the desired setting where all causal predictors are identifiable, and it can happen, for example, when the environments are rich enough. The following proposition shows that $p^{\prime}_{i}$ satisfying (15) is not valid in such settings.
Proposition 3.5.
Let $S^{*}$ be the set of true causal predictors. Consider the desired case where only $H_{0,S^{*}}(\mathcal{E})$ is true, $p_{S^{*}}\sim U(0,1)$ and $p_{S}=0$ for any $S\neq S^{*}$. Then, for $p^{\prime}_{i}$ satisfying (15), there exists some $c^{*}\in(0,1)$ such that $P_{H^{*}_{0,i}}(p^{\prime}_{i}\leq c^{*})>c^{*}$.
This result excludes the possibility of constructing $p_{i}^{\prime}$ based on some analytic function, such that $p_{i}^{\prime}$ is strictly smaller than $p_{i}^{*}$ in probability $1$ (recall that $p_{i}^{*}$ is obtained by a maximum function (7)). Consequently, considering false discovery rate control for ICP does not appear to be a promising idea.
4 Simulations and a real data application
4.1 Numerical simulations
We implement simulations to verify the simultaneous guarantee (13) empirically and to illustrate the extra information one may obtain by using the simultaneous false discovery bound compared to the original ICP approach. All simulations were carried out in R, and the code is available at https://github.com/Jinzhou-Li/ICPsimultaneousBounds.
We generate samples based on the following linear structural equation model:
$X^{e}\leftarrow BX^{e}+\epsilon^{e},$
where $X^{e}=(X^{e}_{1},\dots,X^{e}_{10})^{T}$, $B$ is a strict lower-triangular matrix, $\epsilon^{e}\sim N_{10}(b^{e},\text{diag}(\sigma))$ and $e\in\{1,2,3,4,5\}$. That is, we consider mean-shift intervention and five environments. We treat $X_{6}$ as the response in our simulations.
We implement $500$ simulations. In each simulation, we randomly sample the non-zero elements of matrix $B$ with a non-zero probability $0.8$. Their values are then uniformly sampled from $1$ to $2$. Each entry of $\sigma$ is uniformly sampled from $0.5$ to $1.5$. For the mean-shift intervention, we set $b^{1}=0$. For other $b_{i}$, we randomly select $3$ non-zero entries excluding $X_{6}$. That is, we consider $3$ intervened variables. The non-zero entries of $b_{i}$ are sampled from $N(10,5^{2})$. We generate $100$ samples for each environment and use a significance level $\alpha=0.05$.
For each generated dataset, we apply ICP and record the number of discoveries. Based on the same p-values used by ICP, we calculate the simultaneous false discovery upper bound $t_{\alpha}(R)$ (see (12)) and the true discovery lower bound $|R|-t_{\alpha}$ (see (14)) for all $5122^{9}=$ 512 sets.
Figure 1: The false discovery upper bounds and true discovery lower bounds for all $512$ sets, as well as the size of each set. The first plot is the averaged result over $500$ simulations, and the second and third plots are the result based on two simulated datasets.
Over these $500$ simulations, the empirical probability of $P(|R\cap\mathcal{N}^{*}|\leq t_{\alpha}(R)\text{ for any }R\subseteq[m])$ is $0.988$, which is larger than $0.951-\alpha=$ 0.95, so the simultaneous guarantee (13) empirically holds. The average number of causal predictors is $4.144$, and the average number of discoveries of ICP is $0.52$. The first plot in Figure 1 shows the average false discovery upper bounds and true discovery lower bounds for all $512$ sets, as well as the size of each set. Note that the x-axis is only the index for each set, and we didn’t present which sets these indices refer to for simplicity. Compared to ICP, extra information can be obtained by looking at these bounds. For example, by looking at the true discovery lower bound of the set containing all predictors (with set index $512$), we know that, on average, there are about three true causal predictors in all nine predictors.
To better illustrate the usage of the simultaneous bounds in practice, we look at the results of two simulated datasets rather than the averaged results. These results are shown in the second and third plots of Figure 1, respectively.
For the first simulated dataset, the true causal predictors are $X_{1},X_{2},X_{3}$ and $X_{4}$. ICP discovers $X_{2}$ and $X_{3}$, meaning that with a probability larger than $0.95$, $X_{2}$ and $X_{3}$ are causal predictors. That is all the information we can get by using ICP. By looking at the simultaneous bounds (see the second plot in Figure 1), however, extra information can be obtained. For example, we can see that the sets with indices $131$ and $257$ has true discovery lower bounds $3$ and $4$, respectively. These two sets are $\{X_{1},X_{2},X_{3},X_{4}\}$ and $\{X_{1},X_{2},X_{3},X_{4},X_{5}\}$. Hence, we know that with probability larger than $0.95$, there are at least three causal predictors in $\{X_{1},X_{2},X_{3},X_{4}\}$, and at least four causal predictors in $\{X_{1},X_{2},X_{3},X_{4},X_{5}\}$. Note that the same information that both $X_{2}$ and $X_{3}$ are true causal predictors can also be obtained, because the true discovery lower bound for $\{X_{2},X_{3}\}$ (with set index $13$) is two.
For the second simulated dataset, the true causal predictors are $\{X_{1},X_{2},X_{3},X_{4},X_{5}\}$. ICP returns zero discoveries. In particular, the corresponding p-values (see (7)) for predictors are $1,1,1,1,\\ 0.848,0.329,1,1,1$. Hence, one can not get much information about the causal predictors by using ICP in this case. By looking at the simultaneous bounds (see the third plot in Figure 1), more information can be obtained. For example, we know that with a probability larger than $0.95$, there are at least one causal predictor in $\{X_{3},X_{7}\}$ (with set index $23$), at least two causal predictors in $\{X_{3},X_{5},X_{7},X_{8}\}$ (with set index $164$), at least three causal predictors in $\{X_{1},X_{3},X_{5},X_{7},X_{8},X_{9}\}$ (with set index $406$), and at least four causal predictors in $\{X_{1},X_{2},X_{3},X_{4},X_{5},X_{7},X_{8},X_{9}\}$ (with set index $503$).
4.2 Real data application
We look at a real dataset used in Stock et al., (2003) about the educational attainment of teenagers (Rouse,, 1995). This dataset contains information of $4739$ students from approximately 1100 US high schools, including gender, ethnicity, achievement test score, whether the father or mother is a college graduate, whether the family owns their home, whether the school is in an urban area, county unemployment rate, state hourly wage in manufacturing, average college tuition, whether the family income above 25000 per year, and region.
We follow Peters et al., (2016) to obtain two datasets of two environments, with $2231$ and $2508$ samples, respectively. The response variable $Y$ is a binary variable indicating whether the student attained a BA degree or higher. We use dummy variables to encode the factors in the dataset, resulting in $13$ predictors in total. The goal is to find the causal predictors of the response $Y$.
Figure 2: The false discovery upper bounds, true discovery lower bounds, and the size of set for all $8192$ sets based on the real data.
We first apply ICP with a significance level $\alpha=0.05$, returning one variable ‘score’ as the causal predictor. It seems reasonable that test score has a causal influence on whether the student would attain a BA degree or higher. We then calculate the simultaneous bounds for all $81922^{13}=$ 8192 sets, and the results are shown in Figure 2. More information can be obtained by looking at these bounds. For example, by looking at the set with index $112$, we know that with a probability larger than $0.95$, there are at least two causal predictors in {‘score’, ’fcollege__\__no’, ’mcollege__\__no’}. It seems plausible that the true discovery lower bound is $2$ for this set, because the two variables of whether father or mother went to college are highly correlated, so it is difficult to distinguish between whether both are causal predictors or only one of them is.
Acknowledgement
The author thanks Jelle Goeman and Nicolai Meinshausen for helpful discussions and suggestions on the draft of this paper. The author gratefully acknowledges support by the SNSF Grant P500PT-210978.
Appendix A Supplementary material
A.1 Proof of Proposition 1
Proposition A.1.
For any $c\in(0,1)$, $P_{H^{*}_{0,i}}(p^{*}_{i}\leq c)\leq c$.
Proof.
When $H^{*}_{0,i}$ is true, we have $S^{*}\subseteq[m]\setminus\{i\}$ as $i\not\in S^{*}$, so $p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}\geq p_{S^{*}}$. Therefore, for any $c\in(0,1)$,
$P_{H^{*}_{0,i}}(p^{*}_{i}\leq c)\leq P_{H^{*}_{0,i}}(p_{S^{*}}\leq c)\leq c,$
where the last inequality holds because $H_{0,S^{*}}(\mathcal{E})$ is true. ∎
A.2 Proof of Proposition 2
Proposition A.2.
[…]
To prove Proposition 2, we first introduce the following lemma, which gives the relation between $\widehat{S}(\mathcal{E})=\{i:p^{*}_{i}(\alpha)\leq\alpha\}$ and $\widehat{S}^{\text{ICP}}(\mathcal{E})$.
Lemma A.1.
[…]
We first prove Lemma A.1.
Proof.
If $\widehat{S}(\mathcal{E})=\emptyset$, that is, $p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}>\alpha$ for all $i\in[m]$, then we must have $\widehat{S}^{\text{ICP}}(\mathcal{E})=\emptyset$. Otherwise, if there exists some $j\in\widehat{S}^{\text{ICP}}(\mathcal{E})$, then for any $R$ whose corresponding $p_{R}>\alpha$, we must have $j\in R$ by the definition of $\widehat{S}^{\text{ICP}}(\mathcal{E})$. But $p^{*}_{j}=\max_{S\subseteq[m]\setminus\{j\}}p_{S}>\alpha$ implies that there exists some $R^{\prime}$ not containing $j$ and $p_{R^{\prime}}>\alpha$, which leads to a contradiction. If $\widehat{S}^{\text{ICP}}(\mathcal{E})=\emptyset$, by the assumption in the proposition, there must exist some $R_{1}$ and $R_{2}$ such that $p_{R_{1}}>\alpha$, $p_{R_{2}}>\alpha$ and $R_{1}\cap R_{2}=\emptyset$. Then we must have $\widehat{S}(\mathcal{E})=\emptyset$. Otherwise, if there exists some $j\in\widehat{S}(\mathcal{E})$, that is, $p^{*}_{j}=\max_{S\subseteq[m]\setminus\{j\}}p_{S}\leq\alpha$, then in order for $p_{R_{1}}>\alpha$ and $p_{R_{2}}>\alpha$, both $R_{1}$ and $R_{2}$ must contain $j$, which contradicts the fact that $R_{1}\cap R_{2}=\emptyset$. Now assume both $\widehat{S}(\mathcal{E})$ and $\widehat{S}^{\text{ICP}}(\mathcal{E})$ are non-empty. If $i\in\widehat{S}^{\text{ICP}}(\mathcal{E})$, that is, for any $S$ such that $H_{0,S}(\mathcal{E})$ is not rejected (that is, $p_{S}>\alpha$), we have $i\in S$. So for any $R\subseteq[m]\setminus\{i\}$, $H_{0,R}(\mathcal{E})$ must be rejected, so $p^{*}_{i}=\max_{R\subseteq[m]\setminus\{i\}}p_{R}\leq\alpha$, which implies that $i\in\widehat{S}(\mathcal{E})$. If $i\in\widehat{S}(\mathcal{E})$, we have $p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}\leq\alpha$, which implies that for any $S$ not containing $i$, $H_{0,S}(\mathcal{E})$ must be rejected. So for any $R$ such that $H_{0,R}(\mathcal{E})$ is not rejected, we must have $i\in R$, which implies that $i\in\widehat{S}^{\text{ICP}}(\mathcal{E})$. ∎
Now we prove Proposition 2.
Proof.
If there exists some $R\subseteq[m]$ such that $H_{0,R}(\mathcal{E})$ is not rejected, that is, $p_{R}>\alpha$. Then $\tilde{p}^{*}_{i}(\alpha)=\max\{p^{*}_{i},\Pi_{S\subseteq[m]}\mathbbm{1}_{p_{S}\leq\alpha}\}=\max\{p^{*}_{i},0\}=p^{*}_{i}$, so $\tilde{S}(\mathcal{E})=\widehat{S}(\mathcal{E})=\widehat{S}^{\text{ICP}}(\mathcal{E})$ by Lemma A.1. Now we consider the remaining case that $H_{0,S}(\mathcal{E})$ is rejected for all $S\subseteq[m]$, that is, $p_{S}\leq\alpha$ for all $S\subseteq[m]$. Then, we have $\widehat{S}^{\text{ICP}}(\mathcal{E})=\emptyset$ by definition, and $\tilde{p}^{*}_{i}(\alpha)=\max\{p^{*}_{i},\Pi_{S\subseteq[m]}\mathbbm{1}_{p_{S}\leq\alpha}\}=\max\{p^{*}_{i},1\}=1>\alpha$ for any $i\in[m]$. Hence $\tilde{S}(\mathcal{E})=\emptyset=\widehat{S}^{\text{ICP}}(\mathcal{E})$. ∎
A.3 Proof of Proposition 3
Proposition A.3.
For $\alpha\in(0,1)$, we have $P\left(\widehat{S}(\mathcal{E})\subseteq S^{*}\right)\geq 1-\alpha.$
Proof.
When $H_{0,S^{*}}(\mathcal{E})$ is not rejected, for any $i\not\in S^{*}$, we have $p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}\geq p_{S^{*}}>\alpha$, so $i\not\in\widehat{S}(\mathcal{E})$. Hence, we have $\widehat{S}(\mathcal{E})\subseteq S^{*}$. Therefore,
$P\left(\widehat{S}(\mathcal{E})\subseteq S^{*}\right)\geq P\left(H_{0,S^{*}}(\mathcal{E})\text{ is not rejected}\right)\geq 1-\alpha.$
A.4 Proof of Proposition 4
Proposition A.4.
For any $c\in(0,1)$, $P_{H^{*}_{0,S}}(p^{*}_{S}\leq c)\leq c$.
Proof.
When $H^{*}_{0,S}$ is true, we have $S^{*}\subseteq[m]\setminus S$ as $S\cap S^{*}=\emptyset$, so $p^{*}_{S}=\max_{I\subseteq[m]\setminus S}p_{I}\geq p_{S^{*}}$. Therefore, for any $c\in(0,1)$,
$P_{H^{*}_{0,S}}(p^{*}_{S}\leq c)\leq P_{H^{*}_{0,S}}(p_{S^{*}}\leq c)\leq c.$
A.5 Proof of Theorem 1
Theorem A.1.
Let $\alpha\in(0,1)$, then
$P(|R\cap\mathcal{N}^{*}|\leq t_{\alpha}(R)\text{ for any }R\subseteq[m])\geq 1-\alpha.$
Proof.
For any $R\subseteq[m]$,
$p^{*}_{R\cap\mathcal{N}^{*}}=\max_{I\subseteq[m]\setminus(R\cap\mathcal{N}^{*})}p_{I}\geq\max_{I\subseteq[m]\setminus\mathcal{N}^{*}}p_{I}=p^{*}_{\mathcal{N}^{*}}.$
So if $p^{*}_{\mathcal{N}^{*}}>\alpha$, we have $p^{*}_{R\cap\mathcal{N}^{*}}>\alpha$, which implies that $|R\cap\mathcal{N}^{*}|\leq t_{\alpha}(R)$ by definition. Hence
$P(|R\cap\mathcal{N}^{*}|\leq t_{\alpha}(R)\text{ for any }R\subseteq[m])\geq P(p^{*}_{\mathcal{N}^{*}}>\alpha).$
In addition, we have
$P(p^{*}_{\mathcal{N}^{*}}\leq\alpha)=P(\max_{I\subseteq[m]\setminus\mathcal{N}^{*}}p_{I}\leq\alpha)\leq P(p_{S^{*}}\leq\alpha)\leq\alpha,$
or equivalently
$P(p^{*}_{\mathcal{N}^{*}}>\alpha)\geq 1-\alpha,$
which completes the proof. ∎
A.6 Proof of Proposition 5
Proposition A.5.
Let $S^{*}$ be the set of true causal predictors. Consider the desired case where only $H_{0,S^{*}}(\mathcal{E})$ is true, $p_{S^{*}}\sim U(0,1)$ and $p_{S}=0$ for any $S\neq S^{*}$. Then, for $p^{\prime}_{i}$ satisfying (15), there exists some $c^{*}\in(0,1)$ such that $P_{H^{*}_{0,i}}(p^{\prime}_{i}\leq c^{*})>c^{*}$.
Proof.
For any $i$ such that $H^{*}_{0,i}$ is true, we have $p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}=p_{S^{*}}$. Since $P_{H^{*}_{0,i}}(p^{\prime}_{i}<p^{*}_{i})>0$, there must exist some rational number $c^{*}\in(0,1)$ such that
$P_{H^{*}_{0,i}}(p^{\prime}_{i}\leq c^{*}<p^{*}_{i})>0.$ (16)
Then,
$P_{H^{*}_{0,i}}(p^{\prime}_{i}\leq c^{*})$
$=$ $P_{H^{*}_{0,i}}(p^{\prime}_{i}\leq c^{*},p^{\prime}_{i}<p^{*}_{i})+P_{H^{*}_{0,i}}(p^{\prime}_{i}\leq c^{*},p^{\prime}_{i}=p^{*}_{i})$
$=$ $P_{H^{*}_{0,i}}(p^{\prime}_{i}\leq c^{*},p^{\prime}_{i}<p^{*}_{i}\leq c^{*})+P_{H^{*}_{0,i}}(p^{\prime}_{i}\leq c^{*}<p^{*}_{i})+P_{H^{*}_{0,i}}(p^{\prime}_{i}\leq c^{*},p^{\prime}_{i}=p^{*}_{i})$
$>$ $P_{H^{*}_{0,i}}(p^{*}_{i}\leq c^{*},p^{\prime}_{i}<p^{*}_{i})+P_{H^{*}_{0,i}}(p^{*}_{i}\leq c^{*},p^{\prime}_{i}=p^{*}_{i})$
$=$ $P_{H^{*}_{0,i}}(p^{*}_{i}\leq c^{*})$
$=$ $P_{H^{*}_{0,i}}(p_{S^{*}}\leq c^{*})$
$=$ $c^{*},$
which completes the proof. ∎

### v3

@@
-Simultaneous false discovery bounds for invariant causal prediction
+On the error control of invariant causal prediction
 Abstract
-Invariant causal prediction (ICP, Peters et al., (2016)) provides a novel way to identify causal predictors of a response by utilizing heterogeneous data from different environments.
-One advantage of ICP is that it guarantees to make no false causal discoveries with high probability.
-Such a guarantee, however, can be too conservative in some applications, resulting in few or no discoveries.
-To address this, we propose simultaneous false discovery bounds for ICP, which provides users with extra flexibility in exploring causal predictors and can extract more informative results.
-These additional inferences come for free, in the sense that they do not require additional assumptions, and the same information obtained by the original ICP is retained.
-We demonstrate the practical usage of our method through simulations and a real dataset.
+Invariant causal prediction (ICP, Peters et al., (2016)) provides a novel way for identifying causal predictors of a response by utilizing heterogeneous data from different environments.
+One notable advantage of ICP is that it guarantees to make no false causal discoveries with high probability.
+Such a guarantee, however, can be overly conservative in some applications, resulting in few or no causal discoveries.
+This raises a natural question: Can we use less conservative error control guarantees for ICP so that more causal information can be extracted from data?
+We address this question in the paper.
+We focus on two commonly used and more liberal guarantees: false discovery rate control and simultaneous true discovery bound.
+Unexpectedly, we find that false discovery rate does not seem to be a suitable error criterion for ICP.
+The simultaneous true discovery bound, on the other hand, proves to be an ideal choice, enabling users to explore potential causal predictors and extract more causal information.
+Importantly, the additional information comes for free, in the sense that no extra assumptions are required and the discoveries from the original ICP approach are fully retained.
+We demonstrate the practical utility of our method through simulations and a real dataset about the educational attainment of teenagers in the US.
 1 Introduction
-Discovering causal predictors of the response of interest is usually the primary goal of scientific research.
-Based solely on observational data, the causal relationship might be unidentifiable.
-In such cases, heterogeneous data from different environments can be helpful for further identifiability and thus is a valuable source for causal discovery.
-Under the multi-environments setting, Peters et al., (2016) proposed a novel method, called invariant causal prediction (ICP), to identify causal predictors of a response by exploiting the invariance of the conditional distribution of the response given its causal predictors across environments.
-Compared to other causal discovery algorithms, one main advantage of ICP is that it provides a statistical confidence guarantee for its output $\widehat{S}^{\text{ICP}}$:
-$P\left(\widehat{S}^{\text{ICP}}\subseteq S^{*}\right)\geq 1-\alpha,$
-where $\alpha\in(0,1)$ is a nominal level and $S^{*}$ denotes the set of true causal predictors.
-That is, ICP ensures that all its discoveries are true causal predictors with a probability larger than $1-\alpha$.
-This is also known as the familywise error rate control guarantee.
-The original ICP approach has been generalized to non-linear setting (Heinze-Deml et al.,, 2018), sequential data (Pfister et al.,, 2019), and transformation models (Kook et al.,, 2023).
-All these methods inherit the familywise error rate control guarantee.
-However, controlling the familywise error rate can be too conservative in many applications, especially when causal predictors are highly correlated with non-causal ones.
+Discovering causal predictors of the response of interest is one of the principal goals of scientific research.
+However, based solely on observational data, causal relationships might be unidentifiable.
+In such cases, heterogeneous data from different environments can enhance identifiability and thus is a valuable source for causal discovery.
+In the multi-environment setting, Peters et al., (2016) proposed a novel method called invariant causal prediction (ICP), which identifies causal predictors by exploiting the invariance of the conditional distribution of the response given its causal predictors across environments.
+Specifically, ICP generates p-values $p_{S}$ for each set $S\subseteq[m]=\{1,\dots,m\}$, where $m$ is the number of variables, by testing whether $S$ is invariant across environments (see (1)).
+The estimated set of causal predictors, $\widehat{S}^{\text{ICP}}$ (see (3)), is then obtained based on these p-values.
+One notable advantage of ICP is that it guarantees, with high probability, that all its discoveries are true causal predictors:
+$P\left(\widehat{S}^{\text{ICP}}\subseteq S^{*}\right)=P\left(|\widehat{S}^{\text{ICP}}\cap\mathcal{N}^{*}|=0\right)\geq 1-\alpha,$
+where $\alpha\in(0,1)$ is a nominal level, $S^{*}$ denotes the set of true causal predictors, and $\mathcal{N}^{*}=[m]\setminus S^{*}$ denotes the set of null causal predictors.
+This guarantee is known as familywise error rate (FWER) control.
+Controlling the FWER, however, can be overly conservative in some applications, particularly when causal predictors are highly correlated with non-causal ones.
 In such cases, ICP may result in few or no causal discoveries.
-To address this, Heinze-Deml et al., (2018) proposed the so-called defining sets, which are the smallest sets guaranteeing to contain at least one true causal predictor.
-They are especially useful in the case where ICP returns an empty set.
-In this paper, we propose a method to obtain simultaneous false discovery bound (Genovese and Wasserman,, 2006; Goeman and Solari,, 2011; Goeman et al.,, 2021) for ICP.
-Specifically, let $\mathcal{N}^{*}\subseteq[m]=\{1,\dots,m\}$ be the index set of non-causal predictors among $m$ predictors, the simultaneous false discovery bound is a function $t_{\alpha}:2^{[m]}\rightarrow\{0,\dots,m\}$, where $2^{m}$ denotes the power set of $[m]$, such that
-$P(|R\cap\mathcal{N}^{*}|\leq t_{\alpha}(R)\text{ for any }R\subseteq[m])\geq 1-\alpha.$
-As a result, one can freely check any set $R$ of interest, and a high probability upper bound of the false discoveries can be obtained by $t_{\alpha}(R)$.
-This contains the familywise error rate control and the defining sets of Heinze-Deml et al., (2018) as special cases.
-In particular, by searching for the largest set $R$ whose corresponding $t_{\alpha}(R)$ is 00, we obtain a set with familywise error rate control; by searching for the smallest sets whose corresponding $t_{\alpha}(R)\leq|R|-1$, we obtain all defining sets.
-Extra causal information can be extracted using the simultaneous false discovery bound.
+This raises the question: Is it possible to control more liberal error rates for ICP, thereby extracting more causal information?
+To answer this question, we focus on two commonly used, less conservative error control guarantees: false discovery rate (FDR) control (Benjamini and Hochberg,, 1995; Benjamini and Yekutieli,, 2001) and the simultaneous true discovery bound (Genovese and Wasserman,, 2006; Goeman and Solari,, 2011).
+Formally, based on the p-values $p_{S}$ generated by ICP, we study the following two problems:
+  • 1.
+    FDR control: Can we obtain a discovery set $\widehat{S}$ satisfying the FDR control guarantee
+    $\mathbb{E}\left[\frac{|\widehat{S}\cap\mathcal{N}^{*}|}{\max\{1,|\widehat{S}|\}}\right]\leq\alpha?$
+    In particular, since FWER control implies FDR control (so $\widehat{S}^{\text{ICP}}$ already guarantees FDR control), the goal here is not merely to obtain a set $\widehat{S}$ with FDR control, but also to have $\widehat{S}$ contain more discoveries than $\widehat{S}^{\text{ICP}}$.
+  • 2.
+    Simultaneous true discovery bound: Can we obtain a function $t_{\alpha}:2^{[m]}\rightarrow[0,1]$, where $2^{[m]}$ denotes the power set of $[m]$, such that
+    $P(|R\cap S^{*}|\geq t_{\alpha}(R)\text{ for any }R\subseteq[m])\geq 1-\alpha?$
+    This bound enables users to freely check any set $R$ they consider interesting, which may be formed based on applying some machine learning methods or their expertise, with $t_{\alpha}(R)$ providing a high probability lower bound on the number of true discoveries in $R$.
+To address these two problems, our main idea is to consider a multiple testing problem (see (4)) that directly tests whether causal predictors are null, and propose p-values based on $p_{S}$ generated by ICP.
+We can then apply false discovery rate control procedures or construct simultaneous true discovery bounds using these p-values.
+To this end, we propose valid p-values $p_{i}^{*}$ (see (5)).
+Furthermore, we prove that the original ICP method is equivalent to directly comparing certain p-values (see (6)), which are slightly more conservative than $p_{i}^{*}$, to the significance level.
+For the first problem, our answer is negative.
+We find that directly comparing $p_{i}^{*}$ to the nominal level already controls the FWER, which is somewhat uncommon because multiple testing procedures, such as Bonferroni’s correction, are typically required to control the FWER.
+Given this, there is no reason to further applying FDR control procedures to $p_{i}^{*}$, and the only hope for obtaining more causal discoveries through FDR control would be if $p^{*}_{i}$ were conservative.
+If that were the case, we may propose better p-values and obtain more causal discoveries by applying FDR control methods to them.
+As a result, we investigate whether $p_{i}^{*}$ is conservative in a special setting where the optimal $p_{S}$ can be obtained by ICP.
+We show that $p_{i}^{*}$ is not conservative in this setting, indicating that obtaining smaller valid p-values is unlikely.
+Consequently, we believe that it is unlikely to obtain more causal discoveries by controlling the FDR for ICP.
+For the second problem, our answer is positive.
+The multiple testing framework provides a convenient way to apply the closed testing procedure from Goeman and Solari, (2011) to obtain simultaneous true discovery bounds.
+Furthermore, we show that due to the special construction of $p_{i}^{*}$, the closed testing adjustment step is unnecessary, which significantly reduces computation as we do not need to implement all steps of the standard closed testing procedure.
+We also present a simultaneous bound based on the original ICP formulation, and we prove its equivalence to the one obtained by closed testing.
+This connection is useful as it provides a better understanding of the statistical properties, such as admissibility, of the simultaneous true discovery bound.
+These properties are less evident in the original ICP formulation but have been well-studied in the context of closed testing (Goeman et al.,, 2021).
+The simultaneous true discovery bound preserves the same causal information obtained by the original ICP approach, and it can extract additional causal insights.
+Hence, it is always beneficial to use it in practice.
 As a quick illustration, for a simulated dataset with nine predictors shown in Section 5.1, no information is obtained by the original ICP approach as it returns zero discoveries.
-However, more causal information can be extracted by using our simultaneous false discovery bounds.
+However, more causal information can be extracted by using the simultaneous true discovery bounds.
 For example, we can know that with a probability larger than $0.95$, there is at least one causal predictor in $\{X_{3},X_{7}\}$, at least two causal predictors in $\{X_{3},X_{5},X_{7},X_{8}\}$, and at least four causal predictors in all nine predictors.
-More simulations and a real data application can be found in Section 4.1.
-To obtain the simultaneous false discovery bound for ICP, we proceed by considering a multiple testing problem (see (6)) that directly tests causal predictors.
-We show that the original ICP approach is equivalent to directly comparing certain p-values (see (8)) to the significance level.
-Then, by generalizing the null hypothesis (6) and p-value (7) to (10) and (11), respectively, we propose a simultaneous false discovery upper bound (12).
-This upper bound can be seen as a special case of the closed testing method by Goeman and Solari, (2011), but in our case, no closed testing adjustment is needed due to the specific form of the p-values.
-In comparison to the original ICP approach, our method provides additional causal information at the cost of testing more hypotheses, without introducing any additional statistical assumptions.
-We also discuss the idea of controlling the false discovery rate for ICP.
+More simulations and a real data application can be found in Section 5.
+We note that the simultaneous true discovery bound $t_{\alpha}(R)$ is equivalent to the simultaneous false discovery bound, in the sense that based on $t_{\alpha}(R)$, we can obtain simultaneous false discovery bound $|R|-t_{\alpha}(R)$ satisfying
+$P(|R\cap\mathcal{N}^{*}|\leq|R|-t_{\alpha}(R)\text{ for any }R\subseteq[m])\geq 1-\alpha.$
+We focus on the simultaneous true discovery bound throughout this paper because it offers a more convenient interpretation for applications.
+Many other error rates, such as k-FWER (Lehmann and Romano,, 2005; Sarkar,, 2007) and false discovery exceedance (Dudoit et al.,, 2004; Korn et al.,, 2004; Romano and Shaikh,, 2006), can be seen as special cases of the simultaneous true discovery bound, see Goeman et al., (2021) for further discussions.
+At last, Heinze-Deml et al., (2018) also investigated the question of mitigating the conservativeness of FWER control for ICP, and they proposed the so-called defining sets to tackle this problem.
+Defining sets are the smallest sets guaranteed to contain at least one true causal predictor, making them particularly useful when ICP returns an empty set.
+As we will see later, the true discovery bound encompasses defining sets as special cases.
 2 A brief recap of invariant causal prediction
-For the sake of simplicity, we use linear models (Peters et al.,, 2016) for illustration. Consider the following linear structural equation model:
-$Y^{e}\leftarrow\beta_{1}X^{e}_{1}+\cdots\beta_{m}X^{e}_{m}+\epsilon^{e}=\beta_{S^{*}}^{T}X^{e}_{S^{*}}+\epsilon^{e},$ (1)
+For the sake of simplicity, we use linear models (Peters et al.,, 2016) for illustration.
+Our results can also be applied to non-linear cases (Heinze-Deml et al.,, 2018) directly.
+Consider the following linear structural equation model:
+$Y^{e}\leftarrow\beta_{1}X^{e}_{1}+\cdots\beta_{m}X^{e}_{m}+\epsilon^{e}=\beta_{S^{*}}^{T}X^{e}_{S^{*}}+\epsilon^{e},$
 […]
-The key idea of ICP is to test whether the conditional distribution of $Y^{e}$ is invariant across environments given predictors $X^{e}_{S}$ for some $S\subseteq[m]$.
+The key idea of ICP is to test whether the conditional distribution of $Y^{e}$ is invariant across environments given predictors $X^{e}_{S}$ for each $S\subseteq[m]$.
 This leads to the following null hypothesis:
@@
 […]
-Under the assumptions that there is no latent variable and $\mathcal{E}$ does not contain the environment where $Y$ is intervened, it is clear that $H_{0,S^{*}}(\mathcal{E})$ is true.
+Under the assumptions that there is no latent variables and $\mathcal{E}$ does not contain the environment where $Y$ is intervened, it is clear that $H_{0,S^{*}}(\mathcal{E})$ is true.
 However, there may be other set $S$ than $S^{*}$ for which $H_{0,S}(\mathcal{E})$ is true, causing an identifiability issue.
@@
 Note that $S(\mathcal{E})\subseteq S^{*}$ as $H_{0,S^{*}}(\mathcal{E})$ is true.
-ICP estimates the causal predictors by using the sample version of $S(\mathcal{E})$ in two steps: (i) For every $S\subseteq[m]$, test $H_{0,S}(\mathcal{E})$ at level $\alpha\in(0,1)$.
-(ii) Obtain the selected causal predictors by
+ICP estimates the causal predictors by using the sample version of $S(\mathcal{E})$ in two steps:
+  • (i)
+    For each $S\subseteq[m]$, test $H_{0,S}(\mathcal{E})$ at level $\alpha\in(0,1)$ to obtain p-value $p_{S}$.
+  • (ii)
+    Obtain the estiamted causal predictors by
     $\widehat{S}^{\text{ICP}}(\mathcal{E})=\underset{S:H_{0,S}(\mathcal{E})\text{ is not rejected }}{\bigcap}S.$ (3)
-The familywise error rate guarantee of $\widehat{S}^{\text{ICP}}(\mathcal{E})$ holds at level $\alpha$ because
-$P\left(\widehat{S}^{\text{ICP}}(\mathcal{E})\subseteq S^{*}\right)=P\left(\underset{S:H_{0,S}(\mathcal{E})\text{ is not rejected }}{\bigcap}S\subseteq S^{*}\right)\geq P(H_{0,S^{*}}(\mathcal{E})\text{ is not rejected })\geq 1-\alpha.$ (5)
-3 Simultaneous false discovery bounds for ICP
-3.1 A multiple testing formulation of ICP
-Instead of directly estimating $S(\mathcal{E})$ as in (4), another natural way is to form a multiple testing problem:
-$H^{*}_{0,i}:i\in\mathcal{N}^{*},\quad i\in[m],$ (6)
-where $\mathcal{N}^{*}=[m]\setminus S^{*}$. To this end, we propose the following p-value for $H^{*}_{0,i}$:
-$p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S},$ (7)
-where $p_{S}$ is a valid p-value for testing $H_{0,S}(\mathcal{E})$ (see (2)), that is, $P_{H_{0,S}(\mathcal{E})}(p_{S}\leq c)\leq c$ for any $c\in(0,1)$.
-Note that it can be obtained in the first step of ICP.
-All p-values in this paper depend on the environment set $\mathcal{E}$, but we omit this dependence for simplicity of notation.
-The validity of $p^{*}_{i}$ is shown in the following proposition.
+The FWER guarantee of $\widehat{S}^{\text{ICP}}(\mathcal{E})$ holds at level $\alpha$ because
+$P\left(\widehat{S}^{\text{ICP}}(\mathcal{E})\subseteq S^{*}\right)=P\left(\underset{S:H_{0,S}(\mathcal{E})\text{ is not rejected }}{\bigcap}S\subseteq S^{*}\right)\geq P(H_{0,S^{*}}(\mathcal{E})\text{ is not rejected })\geq 1-\alpha.$
+3 A multiple testing formulation of ICP
+The original formulation of ICP described in Section 2, while elegant and concise, may not be the most suitable for addressing the two problems described in Section 1.
+In this section, we provide a multiple testing perspective on the original ICP approach, which offers additional insights for addressing the considered problems.
+Specifically, instead of directly estimating $S(\mathcal{E})$ as in (3), another natural approach is to consider the following multiple testing problem:
+$H^{*}_{0,i}:i\in\mathcal{N}^{*},\quad i\in[m].$ (4)
+Recall that $\mathcal{N}^{*}=[m]\setminus S^{*}$ is the set of null causal predictors.
+The idea is to obtain valid p-values for these hypotheses, then apply false discovery rate control procedures, such as the Benjamini-Yekutieli procedure (Benjamini and Yekutieli,, 2001), or construct simultaneous true discovery bounds using these p-values.
+These approaches may lead to more discoveries compared to $\widehat{S}^{\text{ICP}}(\mathcal{E})$, because they are less conservative than controlling the FWER.
+To this end, based on the p-values $p_{S}$ generated by the first step of ICP, we propose the following p-value for $H^{*}_{0,i}$:
+$p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}.$ (5)
+All p-values in this paper depend on the environment set $\mathcal{E}$, but we surpress this dependence for simplicity in notation.
+We take the p-values $p_{S}$ as given from the ICP approach and do not elaborate on their calculation.
+For more details, please refer to Peters et al., (2016).
+As shown in the following proposition, $p^{*}_{i}$ is a valid p-value.
 Proposition 3.1.
 For any $c\in(0,1)$, $P_{H^{*}_{0,i}}(p^{*}_{i}\leq c)\leq c$.
-To connect to the original ICP approach, we need to introduce a slightly more complicated p-value. For a given $\tau\in(0,1)$, let
+One may wonder how the original ICP method connects to this multiple testing formulation, and we now establish this connection.
+Since the original ICP method controls the FWER, one might guess that applying some FWER control procedure, such as the Bonferroni correction or Holm’s procedure (Holm,, 1979), would yield a set equivalent to $\widehat{S}^{\text{ICP}}(\mathcal{E})$.
+However, this is not the case.
+Perhaps surprisingly, $\widehat{S}^{\text{ICP}}(\mathcal{E})$ is almost equivalent to the set obtained by directly comparing $p^{*}_{i}$ to the nominal level, without applying any multiple testing corrections.
+We need to introduce a slightly more complicated p-value than $p^{*}_{i}$ in order to establish the connection. For a given $\tau\in(0,1)$, let
 $\tilde{p}^{*}_{i}(\tau)=\max\{p^{*}_{i},\Pi_{S\subseteq[m]}\mathbbm{1}_{p_{S}\leq\tau}\}.$ (6)
-Note that $\tilde{p}^{*}_{i}(\tau)$ is also valid because $\tilde{p}^{*}_{i}(\tau)\geq p^{*}_{i}$.
-Then, as the following proposition shows, the discovery set obtained by using ICP is equivalent to the one obtained by directly comparing $\tilde{p}^{*}_{i}(\alpha)$ to the significance level $\alpha$.
+Note that $\tilde{p}^{*}_{i}(\tau)$ is also a valid p-value because $\tilde{p}^{*}_{i}(\tau)\geq p^{*}_{i}$.
+Then, as the following proposition shows, the discovery set $\widehat{S}^{\text{ICP}}(\mathcal{E})$ obtained by ICP is equivalent to the set obtained by directly comparing $\tilde{p}^{*}_{i}(\alpha)$ to the nominal level $\alpha$.
 Proposition 3.2.
-[…]
-Hence, $\tilde{S}(\mathcal{E})$ possesses the familywise error rate control guarantee (see (5)). In fact, the discovery set
+For $\alpha\in(0,1)$, let $\tilde{S}(\mathcal{E})=\{i:\tilde{p}^{*}_{i}(\alpha)\leq\alpha\}$. Then $\tilde{S}(\mathcal{E})=\widehat{S}^{\text{ICP}}(\mathcal{E})$.
+As a result, $\tilde{S}(\mathcal{E})$ possesses the FWER control guarantee. In fact, the discovery set
 $\widehat{S}(\mathcal{E})=\{i:p^{*}_{i}\leq\alpha\}$ (7)
-obtained by directly comparing $p^{*}_{i}$ to the significance level also controls the familywise error rate, as shown in the following proposition.
+obtained by directly comparing $p^{*}_{i}$ to the nominal level $\alpha$ also controls the FWER, as shown in the following proposition.
 Proposition 3.3.
 For $\alpha\in(0,1)$, we have $P\left(\widehat{S}(\mathcal{E})\subseteq S^{*}\right)\geq 1-\alpha.$
-Therefore, both $\widehat{S}(\mathcal{E})$ and $\widehat{S}^{\text{ICP}}(\mathcal{E})$ (equivalent to $\tilde{S}(\mathcal{E})$) control the familywise error rate, and using $\widehat{S}(\mathcal{E})$ seems better because $p^{*}_{i}\leq\tilde{p}^{*}_{i}(\alpha)$.
-However, the scenario where $p^{*}_{i}<\tilde{p}^{*}_{i}(\alpha)$ only occurs when $p_{S}\leq\alpha$ for all $S\subseteq[m]$.
-In such cases, we have $\widehat{S}^{\text{ICP}}(\mathcal{E})=\emptyset$ and $\widehat{S}(\mathcal{E})=[m]$, which appears not meaningful.
+Note that $\widehat{S}^{\text{ICP}}(\mathcal{E})=\tilde{S}(\mathcal{E})\subseteq\widehat{S}(\mathcal{E})$ as $p^{*}_{i}\leq\tilde{p}^{*}_{i}(\alpha)$.
+Given that both $\widehat{S}(\mathcal{E})$ and $\widehat{S}^{\text{ICP}}(\mathcal{E})$ control the FWER, using $\widehat{S}(\mathcal{E})$ seems to improve upon the original ICP approach.
+However, $p^{*}_{i}<\tilde{p}^{*}_{i}(\alpha)$ only occurs when $p_{S}\leq\alpha$ for all $S\subseteq[m]$.
+In such cases, $\widehat{S}^{\text{ICP}}(\mathcal{E})=\emptyset$ and $\widehat{S}(\mathcal{E})=[m]$, which seems useless because even $p_{S^{*}}\leq\alpha$, leading to a false rejection of $H_{0,S^{*}}(\mathcal{E})$.
 In practice, $\widehat{S}(\mathcal{E})$ and $\widehat{S}^{\text{ICP}}(\mathcal{E})$ are generally equivalent.
-3.2 Simultaneous false discovery bounds for ICP
-Based on the multiple testing formulation (6), we propose simultaneous false bounds by considering all intersection hypotheses.
-Specifically, for a given $S\subseteq[m]$, by generalizing (6) and (7), we consider null hypothesis
+4 Can we control less conservative errors for ICP?
+4.1 FDR control does not seem well-suited for ICP
+Proposition 3.3 shows that directly comparing $p^{*}_{1},\dots,p^{*}_{m}$ to the nominal level automatically ensures FWER control, as if there were no multiple testing issue.
+This is somewhat unexpected, as one would typically need to apply multiple testing adjustments (e.g., Bonferroni correction) to control the FWER.
+Consequently, there is no need to further apply FDR control procedures to $p^{*}_{1},\dots,p^{*}_{m}$.
+Given this, the only remaining hope for obtaining more causal discoveries through FDR control is that the construction (5) of $p^{*}_{i}$ based on $p_{S}$ is conservative.
+Then, if a more powerful method for obtaining smaller p-values for $H^{*}_{0,i}$ based on $p_{S}$ can be proposed, we might return to the “normal” situation where directly comparing these p-values to the nominal level no longer controls the FWER.
+If this could be achieved, applying FDR control procedures may lead to more causal discoveries.
+Unfortunately, as we will show below, $p^{*}_{i}$ defined by (5) does not appear to be conservative.
+We consider a special case where the optimal $p_{S}$ can be obtained for testing $H_{0,S}(\mathcal{E})$.
+That is, we have $p_{S}\sim U(0,1)$ for true null hypothesis and $p_{S}=0$ for false null hypothesis.
+This represents the most desirable scenario, where powerful tests for $H_{0,S}(\mathcal{E})$ are available.
+It is also called the Dirac-uniform configuration (Finner et al.,, 2009).
+Note that obtaining these $p_{S}$ values is a separate task from constructing a valid p-value for the hypothesis (4) based on them.
+Then, for $i\in S(\mathcal{E})$ (see (2)), which is an identifiable causal predictor based on ICP, the construction (5) yields
+$p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}=0,$
+because $H_{0,S}(\mathcal{E})$ is false for any set $S$ that does not contain $i$, implying that $p_{S}=0$.
+In other words, for any identifiable causal predictor, the corresponding p-value will be 00, which is the best possible outcome.
+Therefore, (5) does not seem to be a conservative construction for $p^{*}_{i}$.
+Furthermore, if we are willing to additionally assume the following weak Assumption 4.1, then in the case where the environment set is sufficiently rich, we can show that $p^{*}_{i}$ is admissible, as demonstrated below.
+Assumption 4.1 (Common null events).
+For every measurable event $E$ and every possible data distribution $P_{1}$ and $P_{2}$, if $P_{1}(E)=0$, then $P_{2}(E)=0$.
+This assumption means that the collection of null events is common across all considered possible distributions.
+It is satisfied by most models, provided that some deterministic corner cases are excluded.
+For more discussion, see the Appendix A of Goeman et al., (2021).
+In addition, assume that we are in a scenario where the only true null hypothesis is $H_{0,S^{*}}(\mathcal{E})$.
+This is the desirable setting in which all causal predictors are identifiable, and it can occur, for example, when the environment set is sufficiently rich.
+In this scenario, for any $i$ such that $H^{*}_{0,i}$ is true, we have
+$p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}=p_{S^{*}}.$
+In this special case, all null p-values are equal to $p_{S^{*}}$, providing a scenario where it is clear that directly comparing the p-values to the nominal level controls the FWER. Then, we have
+$P_{H^{*}_{0,i}}(p^{*}_{i}\leq\alpha)=P_{H^{*}_{0,i}}(p_{S^{*}}\leq\alpha)=\alpha.$
+By the following Lemma 4.1, $p^{*}_{i}$ is admissible.
+Note that both obtaining optimal $p_{S}$ and having a sufficiently rich environment set $\mathcal{E}$ are independent of the underlying truth regarding which variables are true causal predictors.
+The admissibility $p^{*}_{i}$ suggests that its construction does not seem to be conservative.
+Lemma 4.1 (Lemma $7$ in the Appendix A of Goeman et al., (2021)).
+Under Assumption 4.1, a p-value $p$ of hypothesis $H$ is admissible if there exist a distribution $P$ under $H$ such that $P_{H}(p\leq\alpha)=\alpha$.
+In general settings where $p_{S}$ is not optimal, we can not fully exclude the possibility of constructing a more powerful (i.e., smaller) $p^{*}_{i}$.
+However, based on the discussions above, we believe it is unlikely that $p^{*}_{i}$ can be significantly improved to yield more discoveries after applying FDR control procedures to the new p-values.
+Therefore, our response to the first question raised in the introduction is negative: FDR control does not seem well-suited for ICP, and the idea of obtaining more discoveries by controlling the FDR in the context of ICP appears unpromising.
+4.2 Simultaneous true discovery bound for ICP
+Now, we address the second question about simultaneous true discovery bound.
+Fortunately, this time, the answer is positive: we can obtain a simultaneous true discovery bound and use it to extract more causal information.
+Below, we present how to obtain this bound.
+Based on the multiple testing formulation (4), a straightforward way to obtain simultaneous true discovery bound is to apply the closed testing procedure (Marcus et al.,, 1976; Goeman and Solari,, 2011) (see Appendix A for a brief recap).
+This requires testing the intersection hypothesis
 $H^{*}_{0,S}:i\in\mathcal{N}^{*}\text{ for all }i\in S$ (8)
-and p-value
+for each $S\subseteq[m]$. To this end, we generalize the p-value (5) and propose the following p-value for testing $H^{*}_{0,S}$:
 $p^{*}_{S}=\max_{I\subseteq[m]\setminus S}p_{I}.$ (9)
-By using a similar argument as in Proposition 3.1, one can see that $p^{*}_{S}$ is valid.
+The validity of $p^{*}_{S}$ is shown in the following proposition.
 Proposition 4.1.
 For any $c\in(0,1)$, $P_{H^{*}_{0,S}}(p^{*}_{S}\leq c)\leq c$.
-For any set $R\subseteq[m]$, let $t_{\alpha}(R)$ be the size of the largest subset of $R$ whose corresponding p-value is larger than $\alpha$. That is,
-$t_{\alpha}(R)=\max\{|I|:I\subseteq R\text{ and }p^{*}_{I}>\alpha\}.$ (12)
-Then, $t_{\alpha}(R)$ is a simultaneous false discovery upper bound, as shown in the following Theorem.
+With valid p-values $p^{*}_{S}$ in hand, we can implement the closed testing procedure to obtain simultaneous true discovery bounds.
+A key observation here is that due to the special construction of $p^{*}_{S}$, there is no need to implement closed testing adjustments.
+To see this, if $p^{*}_{S}\leq\alpha$, we must have $p^{*}_{S’}\leq\alpha$ for all $S\subseteq S’$ ’, indicating that if a hypothesis is locally rejected, it must be rejected by closed testing.
+On the other hand, by definition, if a hypothesis is rejected by closed testing, it must be locally rejected.
+Therefore, a hypothesis is rejected by closed testing if and only if it is locally rejected, hence no closed testing adjustment is needed in this case.
+This leads to the following simultaneous true discovery bound:
+$t_{\alpha}(R)=\min\{|R\setminus I|:I\subseteq R\text{ and }p^{*}_{I}>\alpha\}.$ (10)
+In words, $t_{\alpha}(R)$ is the size of the smallest subset of $R$ whose corresponding complement in $R$ has a p-value larger than $\alpha$.
+It can be seen as a shortcut of the standard closed testing procedure, which simplifies the computation by not requiring the full implementation of all the steps in closed testing.
+The theorem below formally shows that $t_{\alpha}(R)$ is a simultaneous true discovery bound.
 Theorem 4.1.
 Let $\alpha\in(0,1)$, then
-$P(|R\cap\mathcal{N}^{*}|\leq t_{\alpha}(R)\text{ for any }R\subseteq[m])\geq 1-\alpha.$ (13)
-Based on the simultaneous guarantee (13), one can freely check any set $R$ of interest, and a high probability upper bound (12) on the false discoveries can be calculated.
-Equivalently, $|R|-t_{\alpha}(R)$ is a simultaneous true discovery lower bound:
-$P(|R\cap S^{*}|\geq|R|-t_{\alpha}(R)\text{ for any }R\subseteq[m])\geq 1-\alpha.$ (14)
-The simultaneous upper bound (12) is the same as the upper bound proposed by Goeman and Solari, (2011) for closed testing (Marcus et al.,, 1976).
-In fact, our method can be seen as a special case of closed testing.
-Compared to the standard closed testing procedure, the difference is that no closed testing adjustments are needed in our case.
-This is due to the specific form of $p^{*}_{S}$.
-In particular, if $p^{*}_{S}\leq\alpha$, we must have $p^{*}_{S^{\prime}}\leq\alpha$ for all $S\subseteq S^{\prime}$.
-Thus, if a hypothesis is locally rejected, it must be rejected by closed testing, so there is no need to implement closed testing adjustments.
-At last, we mention that calculating the false discovery bound (12) can be computationally intractable for large $|R|$ (or large $m$).
-In fact, computation is also a fundamental issue for the original ICP approach, which requires testing $2^{|m|}$ hypotheses.
-Some tricks are suggested to deal with this issue, including first reducing the size of the potential causal predictors by using a pre-screening method such as Lasso.
-See Peters et al., (2016) for more discussions.
-3.3 A discussion about false discovery rate control for ICP
-With a multiple testing formulation (6) and valid p-values (7), an alternative strategy to obtain a less conservative result than $\widehat{S}^{\text{ICP}}(\mathcal{E})$ is to control the commonly used false discovery rate (Benjamini and Hochberg,, 1995; Benjamini and Yekutieli,, 2001; Lehmann and Romano,, 2022).
-Due to Proposition 3.3, however, directly comparing $p^{*}_{1},\dots,p^{*}_{m}$ to the significance level already controls the familywise error rate, so there is no point in further applying false discovery rate control procedures on these p-values.
-This observation is interesting because, typically, multiple testing corrections (e.g., Bonferroni correction) on raw p-values are required to control the familywise error rate.
-Thus, it is natural to ask: Can we obtain smaller valid p-values than $p^{*}_{1},\dots,p^{*}_{m}$?
-If so, one may apply false discovery rate control procedures to obtain more discoveries than ICP.
-However, in the following, we argue that this approach does not seem promising.
-Specifically, we show that if some $p^{\prime}_{i}$ satisfying
-${\mathbb{P}}(p^{\prime}_{i}\leq p^{*}_{i})=1\quad\text{and}\quad{\mathbb{P}}_{H^{*}_{0,i}}(p^{\prime}_{i}<p^{*}_{i})>0,$ (15)
-then it is not valid in general.
-In particular, suppose that we have an ideal test for $H_{0,S}(\mathcal{E})$ that yields p-values $p_{S}\sim U(0,1)$ for true nulls and $p_{S}=0$ for false nulls.
-Consider a setting where the only true null hypothesis is $H_{0,S^{*}}(\mathcal{E})$.
-This is the desired setting where all causal predictors are identifiable, and it can happen, for example, when the environments are rich enough.
-The following proposition shows that $p^{\prime}_{i}$ satisfying (15) is not valid in such settings.
-Proposition 3.5.
-Let $S^{*}$ be the set of true causal predictors.
-Consider the desired case where only $H_{0,S^{*}}(\mathcal{E})$ is true, $p_{S^{*}}\sim U(0,1)$ and $p_{S}=0$ for any $S\neq S^{*}$.
-Then, for $p^{\prime}_{i}$ satisfying (15), there exists some $c^{*}\in(0,1)$ such that $P_{H^{*}_{0,i}}(p^{\prime}_{i}\leq c^{*})>c^{*}$.
-This result excludes the possibility of constructing $p_{i}^{\prime}$ based on some analytic function, such that $p_{i}^{\prime}$ is strictly smaller than $p_{i}^{*}$ in probability $1$ (recall that $p_{i}^{*}$ is obtained by a maximum function (7)).
-Consequently, considering false discovery rate control for ICP does not appear to be a promising idea.
+$P(|R\cap S^{*}|\geq t_{\alpha}(R)\text{ for any }R\subseteq[m])\geq 1-\alpha.$ (11)
+Based on the simultaneous guarantee (11), users can freely check any set $R$ they regard as interesting, and calculate a high-probability lower bound $t_{\alpha}(R)$ on the true discoveries.
+The simultaneous true discovery bound encompasses both FWER control and the defining sets (Heinze-Deml et al.,, 2018) as special cases.
+Specifically, by searching for the largest set $R$ whose $t_{\alpha}(R)=|R|$, we obtain a set with FWER control; by searching for all the smallest sets $R$ with $t_{\alpha}(R)=1$, we obtain the defining sets.
+In particular, it is easy to check that for $\widehat{S}^{\text{ICP}}$, we have $t_{\alpha}(\widehat{S}^{\text{ICP}})=|\widehat{S}^{\text{ICP}}|$, thereby recovering the FWER control guarantee of ICP.
+Hence, the same causal discoveries and information obtained from the original ICP approach are fully retained.
+This indicates that using our proposed simultaneous true discovery bound to check more sets is always beneficial, which will only extract additional useful information from data.
+See Section 5 for illustrations.
+In fact, one can also obtain a simultaneous true discovery bound based on the original ICP formulation, without resorting to the multiple testing framework  [^1] We thank an anonymous referee for pointing out this..
+Specifically, let
+$\tilde{t}_{\alpha}(R)=\underset{S\subseteq[m]\text{ with }p_{S}>\alpha}{\min}|R\cap S|,\quad\text{for all }R\subseteq[p].$ (12)
+Then, we have
+$P(|R\cap S^{*}|\geq\tilde{t}_{\alpha}(R)\text{ for any }R\subseteq[m])\geq P(H_{0,S^{*}}(\mathcal{E})\text{ is not rejected})\geq 1-\alpha,$
+where the first inequality holds by the definition of $\tilde{t}_{\alpha}(R)$.
+Although obtained based on different ideas, $t_{\alpha}(R)$ and $\tilde{t}_{\alpha}(R)$ (see (10) and (12)) are actually equivalent, as shown in the following theorem.
+Theorem 4.2.
+Let $\alpha\in(0,1)$. For any $R\subseteq[m]$,
+$t_{\alpha}(R)=\tilde{t}_{\alpha}(R).$
+It is not uncommon in the literature that two seemingly different methods are actually equivalent.
+For example, the simultaneous bound based on closed testing from Goeman and Solari, (2011) is equivalent to the bound from Genovese and Wasserman, (2006), which does not use closed testing (Goeman et al.,, 2021).
+Recognizing such connections is valuable because it provides different perspectives on the same method and allows insights to be gained from different angles.
+For example, in our case, understanding that $\tilde{t}_{\alpha}(R)$ is equivalent to a closed testing bound helps clarify its admissibility (Goeman et al.,, 2021), which is rather unclear from its original formulation (12).
 5 Simulations and a real data application
 5.1 Numerical simulations
-We implement simulations to verify the simultaneous guarantee (13) empirically and to illustrate the extra information one may obtain by using the simultaneous false discovery bound compared to the original ICP approach.
+We implement simulations to verify the simultaneous guarantee (11) and to illustrate the extra information one may obtain by using the simultaneous true discovery bound, compared to the original ICP approach.
 All simulations were carried out in R, and the code is available at https://github.com/Jinzhou-Li/ICPsimultaneousBounds.
 We generate samples based on the following linear structural equation model:
-$X^{e}\leftarrow BX^{e}+\epsilon^{e},$
-where $X^{e}=(X^{e}_{1},\dots,X^{e}_{10})^{T}$, $B$ is a strict lower-triangular matrix, $\epsilon^{e}\sim N_{10}(b^{e},\text{diag}(\sigma))$ and $e\in\{1,2,3,4,5\}$.
+$X^{e}\leftarrow BX^{e}+\epsilon^{e}+b^{e},$
+where $X^{e}=(X^{e}_{1},\dots,X^{e}_{10})^{T}$, $B$ is a strict lower-triangular matrix, $\epsilon^{e}\sim N_{10}(0,\text{diag}(\sigma))$ and $e\in\{1,2,3,4,5\}$.
 That is, we consider mean-shift intervention and five environments.
@@
 In each simulation, we randomly sample the non-zero elements of matrix $B$ with a non-zero probability $0.8$.
-Their values are then uniformly sampled from $1$ to $2$.
-Each entry of $\sigma$ is uniformly sampled from $0.5$ to $1.5$.
+Their values are uniformly sampled from $(1,2)$.
+Each entry of $\sigma$ is uniformly sampled from $(0.5,1.5)$.
 For the mean-shift intervention, we set $b^{1}=0$.
-For other $b_{i}$, we randomly select $3$ non-zero entries excluding $X_{6}$.
+For other $b^{i}$, we randomly select $3$ non-zero entries excluding $X_{6}$.
 That is, we consider $3$ intervened variables.
-The non-zero entries of $b_{i}$ are sampled from $N(10,5^{2})$.
+The non-zero entries of $b^{i}$ are sampled from $N(10,5^{2})$.
 We generate $100$ samples for each environment and use a significance level $\alpha=0.05$.
 For each generated dataset, we apply ICP and record the number of discoveries.
-Based on the same p-values used by ICP, we calculate the simultaneous false discovery upper bound $t_{\alpha}(R)$ (see (12)) and the true discovery lower bound $|R|-t_{\alpha}$ (see (14)) for all $5122^{9}=$ 512 sets.
+Based on the same p-values used by ICP, we calculate the simultaneous true discovery upper bound $t_{\alpha}(R)$ (see (10)) for all $5122^{9}=$ 512 sets.
+We also present the false discovery lower bounds $|R|-t_{\alpha}$ for illustration purposes, which essentially convey the same information as the simultaneous true discovery bound.
 Figure 1: The false discovery upper bounds and true discovery lower bounds for all $512$ sets, as well as the size of each set.
 The first plot is the averaged result over $500$ simulations, and the second and third plots are the result based on two simulated datasets.
-Over these $500$ simulations, the empirical probability of $P(|R\cap\mathcal{N}^{*}|\leq t_{\alpha}(R)\text{ for any }R\subseteq[m])$ is $0.988$, which is larger than $0.951-\alpha=$ 0.95, so the simultaneous guarantee (13) empirically holds.
+Over these $500$ simulations, the empirical probability of $P(|R\cap S^{*}|\geq t_{\alpha}(R)\text{ for any }R\subseteq[m])$ is $0.988$, which is larger than $0.951-\alpha=$ 0.95, so the simultaneous guarantee (11) holds.
 The average number of causal predictors is $4.144$, and the average number of discoveries of ICP is $0.52$.
@@
 For example, by looking at the true discovery lower bound of the set containing all predictors (with set index $512$), we know that, on average, there are about three true causal predictors in all nine predictors.
-To better illustrate the usage of the simultaneous bounds in practice, we look at the results of two simulated datasets rather than the averaged results.
+To better illustrate the usage of the simultaneous bounds in practice, we look at two simulated datasets rather than the averaged results.
 These results are shown in the second and third plots of Figure 1, respectively.
-For the first simulated dataset, the true causal predictors are $X_{1},X_{2},X_{3}$ and $X_{4}$.
+For the first simulated dataset, the true causal predictors are $\{X_{1},X_{2},X_{3},X_{4}\}$.
 ICP discovers $X_{2}$ and $X_{3}$, meaning that with a probability larger than $0.95$, $X_{2}$ and $X_{3}$ are causal predictors.
@@
 5.2 Real data application
-We look at a real dataset used in Stock et al., (2003) about the educational attainment of teenagers (Rouse,, 1995).
+We look at a real dataset used in Stock et al., (2003) about the educational attainment of teenagers in the US (Rouse,, 1995).
 This dataset contains information of $4739$ students from approximately 1100 US high schools, including gender, ethnicity, achievement test score, whether the father or mother is a college graduate, whether the family owns their home, whether the school is in an urban area, county unemployment rate, state hourly wage in manufacturing, average college tuition, whether the family income above 25000 per year, and region.
@@
 We first apply ICP with a significance level $\alpha=0.05$, returning one variable ‘score’ as the causal predictor.
-It seems reasonable that test score has a causal influence on whether the student would attain a BA degree or higher.
+It is reasonable that test score has a causal influence on whether the student would attain a BA degree or higher.
 We then calculate the simultaneous bounds for all $81922^{13}=$ 8192 sets, and the results are shown in Figure 2.
@@
 It seems plausible that the true discovery lower bound is $2$ for this set, because the two variables of whether father or mother went to college are highly correlated, so it is difficult to distinguish between whether both are causal predictors or only one of them is.
+6 Discussion
+Motivated by the issue that FWER control in the original ICP approach can be overly conservative for some applications, resulting in few or no causal discoveries and the loss of valuable causal information, we address the error control problem of ICP in this paper.
+We focus on two commonly used and more liberal error control guarantees: FDR control and simultaneous true discovery bounds.
+We investigate whether it is possible to generalize the original ICP approach to possess these two guarantees, thereby revealing more causal information.
+The main idea is to reformulate ICP within the framework of multiple testing.
+We propose valid p-values for this multiple testing problem and show the equivalence of this formulation to the original ICP approach.
+Our answer for FDR control problem is negative: in this special multiple testing problem where the p-values are constructed based on ICP, FDR seems not a suitable error criterion.
+However, for simultaneous true discovery bounds, we provide a positive answer and propose such bounds based on closed testing.
+Through simulations and a real dataset, we demonstrate that using simultaneous true discovery bounds can indeed uncover more valuable causal information compared to the original ICP approach.
+Importantly, this additional information comes at no cost, as the original information is fully retained, making it always advantageous to use in practice.
+The original ICP method has been generalized to various settings, including non-linear models (Heinze-Deml et al.,, 2018), sequential data (Pfister et al.,, 2019), and transformation models (Kook et al.,, 2023).
+It would be interesting to either directly apply or generalize our proposed simultaneous true discovery bound in these settings.
+When testing the null hypothesis $H^{*}_{0,i}$ (see (4)), p-values based on the ICP approach are constructed by testing for invariance across different environments.
+However, other sources of information, such as testing for conditional independence, may also provide evidence to reject $H^{*}_{0,i}$.
+If new p-values can be obtained by combining information from different principles when testing $H^{*}_{0,i}$, then the simultaneous false discovery bound based on closed testing can be directly applied to provide causal discoveries with error control guarantees, and FDR control may become suitable again.
+Note that it is unclear how to achieve this with the original ICP formulation.
+This is another advantage of the multiple testing formulation (4).
+For future research, it would be interesting to develop concrete methods for combining information based on different principles when testing $H^{*}_{0,i}$ to achieve error-controlled causal discoveries.
+Finally, we note that calculating the simultaneous true discovery bound (10) can become computationally intractable for large $|R|$ (or large $m$).
+This computational challenge is also inherent in the original ICP approach, which requires testing $2^{|m|}$ hypotheses.
+Various strategies have been proposed to address this issue, including reducing the set of potential causal predictors through pre-screening methods such as Lasso (see Peters et al., (2016) for further discussion).
+Addressing this computational challenge is another interesting direction for future research.
 Acknowledgement
-The author thanks Jelle Goeman and Nicolai Meinshausen for helpful discussions and suggestions on the draft of this paper. The author gratefully acknowledges support by the SNSF Grant P500PT-210978.
-Appendix A Supplementary material
-A.1 Proof of Proposition 1
-Proposition A.1.
-For any $c\in(0,1)$, $P_{H^{*}_{0,i}}(p^{*}_{i}\leq c)\leq c$.
+Jinzhou Li thanks Nicolai Meinshausen for helpful discussions and suggestions on the draft of this paper. Jinzhou Li gratefully acknowledges support from the SNSF Grant P500PT-210978.
+SUPPLEMENTARY MATERIAL
+Appendix A A brief recap on closed testing
+The closed testing procedure was originally introduced for controlling the familywise error rate (Marcus et al.,, 1976).
+Later, Goeman and Solari, (2011) proposed simultaneous false discovery bounds based on it.
+Below, we present the three main steps of the closed testing procedure.
+Consider a multiple testing problem with hypotheses $H_{1},\dots,H_{m}$.
+For any $I\subseteq[m]$, let $H_{I}=\cap_{i\in I}H_{i}$ be an intersection hypothesis, where $H_{I}$ is true if and only if $H_{i}$ is true for all $i\in I$.
+The closed testing procedure proceeds as follows:
+  • 1.
+    Obtain local test results: Locally test all intersection hypotheses $H_{I}$ for $I\subseteq[m]$ at level $\alpha\in(0,1)$.
+  • 2.
+    Implement the closed testing adjustment: $H_{I}$ is rejected by closed testing if and only if $H_{J}$ is locally rejected in Step 1 for all $J\supseteq I$.
+  • 3.
+    Obtain simultaneous false discovery bound $u^{ct}(R)$: For any set $R\subseteq[m]$, let $u^{ct}(R)$ be the size of the largest subset of $R$ that is not rejected by closed testing in Step 2.
+Then, Goeman and Solari, (2011) proved that $u^{ct}(R)$ is a simultaneous false discovery bound. That is,
+${\mathbb{P}}(|R\cap\mathcal{N}|\leq u^{ct}(R),\forall R\subseteq[m])\geq 1-\alpha,$
+where $\mathcal{N}\subseteq[m]$ is the set of null hypotheses.
+Equivalently, $t^{ct}(R)=|R|-u^{ct}(R)$ is a simultaneous true discovery bound satisfying
+${\mathbb{P}}(|R\cap S^{*}|\geq t^{ct}(R),\forall R\subseteq[m])\geq 1-\alpha,$
+where $S^{*}=[m]\setminus\mathcal{N}$ is the set of true hypotheses.
+In particular, by the definition of $u^{ct}(R)$, we have that $t^{ct}(R)$ is the smallest subset of $R$ whose complement in $R$ is not rejected by closed testing.
+Appendix B Proofs
+B.1 Proof of Proposition 3.1
 Proof.
-When $H^{*}_{0,i}$ is true, we have $S^{*}\subseteq[m]\setminus\{i\}$ as $i\not\in S^{*}$, so $p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}\geq p_{S^{*}}$. Therefore, for any $c\in(0,1)$,
+(Proposition 3.1) When $H^{*}_{0,i}$ is true, we have $S^{*}\subseteq[m]\setminus\{i\}$ as $i\not\in S^{*}$, so $p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}\geq p_{S^{*}}$.
+Therefore, for any $c\in(0,1)$,
 $P_{H^{*}_{0,i}}(p^{*}_{i}\leq c)\leq P_{H^{*}_{0,i}}(p_{S^{*}}\leq c)\leq c,$
 where the last inequality holds because $H_{0,S^{*}}(\mathcal{E})$ is true. ∎
-A.2 Proof of Proposition 2
-Proposition A.2.
-[…]
-To prove Proposition 2, we first introduce the following lemma, which gives the relation between $\widehat{S}(\mathcal{E})=\{i:p^{*}_{i}(\alpha)\leq\alpha\}$ and $\widehat{S}^{\text{ICP}}(\mathcal{E})$.
-Lemma A.1.
-[…]
-We first prove Lemma A.1.
+B.2 Proof of Proposition 3.2
+To prove Proposition 3.2, we first introduce the following lemma, which gives the relation between $\widehat{S}(\mathcal{E})=\{i:p^{*}_{i}\leq\alpha\}$ and $\widehat{S}^{\text{ICP}}(\mathcal{E})$.
+Lemma B.1.
+Assume that there exists $R\subseteq[m]$ such that $H_{0,R}(\mathcal{E})$ is not rejected.
+Let $\widehat{S}(\mathcal{E})=\{i:p^{*}_{i}\leq\alpha\}$.
+Then, $\widehat{S}(\mathcal{E})=\widehat{S}^{\text{ICP}}(\mathcal{E})$.
+We first prove Lemma B.1.
 Proof.
-If $\widehat{S}(\mathcal{E})=\emptyset$, that is, $p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}>\alpha$ for all $i\in[m]$, then we must have $\widehat{S}^{\text{ICP}}(\mathcal{E})=\emptyset$.
+(Lemma B.1) If $\widehat{S}(\mathcal{E})=\emptyset$, that is, $p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}>\alpha$ for all $i\in[m]$, then we must have $\widehat{S}^{\text{ICP}}(\mathcal{E})=\emptyset$.
 Otherwise, if there exists some $j\in\widehat{S}^{\text{ICP}}(\mathcal{E})$, then for any $R$ whose corresponding $p_{R}>\alpha$, we must have $j\in R$ by the definition of $\widehat{S}^{\text{ICP}}(\mathcal{E})$.
 But $p^{*}_{j}=\max_{S\subseteq[m]\setminus\{j\}}p_{S}>\alpha$ implies that there exists some $R^{\prime}$ not containing $j$ and $p_{R^{\prime}}>\alpha$, which leads to a contradiction.
-If $\widehat{S}^{\text{ICP}}(\mathcal{E})=\emptyset$, by the assumption in the proposition, there must exist some $R_{1}$ and $R_{2}$ such that $p_{R_{1}}>\alpha$, $p_{R_{2}}>\alpha$ and $R_{1}\cap R_{2}=\emptyset$.
+If $\widehat{S}^{\text{ICP}}(\mathcal{E})=\emptyset$, by the assumption that there exists $R\subseteq[m]$ such that $H_{0,R}(\mathcal{E})$ is not rejected (i.e., $p_{R}>\alpha$), there are two possibilities: (1) $R=\emptyset$ is the only set whose corresponding $p_{R}>\alpha$.
+In this case, we have $p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}\geq p_{R}>\alpha$ for all $i\in[m]$, so $\widehat{S}(\mathcal{E})=\emptyset$.
+(2) There exist two different sets $R_{1}$ and $R_{2}$ such that $p_{R_{1}}>\alpha$, $p_{R_{2}}>\alpha$ and $R_{1}\cap R_{2}=\emptyset$.
 Then we must have $\widehat{S}(\mathcal{E})=\emptyset$.
 Otherwise, if there exists some $j\in\widehat{S}(\mathcal{E})$, that is, $p^{*}_{j}=\max_{S\subseteq[m]\setminus\{j\}}p_{S}\leq\alpha$, then in order for $p_{R_{1}}>\alpha$ and $p_{R_{2}}>\alpha$, both $R_{1}$ and $R_{2}$ must contain $j$, which contradicts the fact that $R_{1}\cap R_{2}=\emptyset$.
+Hence, $\widehat{S}^{\text{ICP}}(\mathcal{E})=\emptyset$ if and only if $\widehat{S}(\mathcal{E})=\emptyset$.
 Now assume both $\widehat{S}(\mathcal{E})$ and $\widehat{S}^{\text{ICP}}(\mathcal{E})$ are non-empty.
-If $i\in\widehat{S}^{\text{ICP}}(\mathcal{E})$, that is, for any $S$ such that $H_{0,S}(\mathcal{E})$ is not rejected (that is, $p_{S}>\alpha$), we have $i\in S$.
+If $i\in\widehat{S}^{\text{ICP}}(\mathcal{E})$, that is, for any $S$ such that $H_{0,S}(\mathcal{E})$ is not rejected, we have $i\in S$.
 So for any $R\subseteq[m]\setminus\{i\}$, $H_{0,R}(\mathcal{E})$ must be rejected, so $p^{*}_{i}=\max_{R\subseteq[m]\setminus\{i\}}p_{R}\leq\alpha$, which implies that $i\in\widehat{S}(\mathcal{E})$.
 If $i\in\widehat{S}(\mathcal{E})$, we have $p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}\leq\alpha$, which implies that for any $S$ not containing $i$, $H_{0,S}(\mathcal{E})$ must be rejected.
-So for any $R$ such that $H_{0,R}(\mathcal{E})$ is not rejected, we must have $i\in R$, which implies that $i\in\widehat{S}^{\text{ICP}}(\mathcal{E})$.
-∎
-Now we prove Proposition 2.
+So for any $R$ such that $H_{0,R}(\mathcal{E})$ is not rejected, we must have $i\in R$.
+Combining with the assumption that there exists $R\subseteq[m]$ such that $H_{0,R}(\mathcal{E})$ is not rejected, we have $i\in\widehat{S}^{\text{ICP}}(\mathcal{E})$.
+Combining the above cases, we have $\widehat{S}^{\text{ICP}}(\mathcal{E})=\widehat{S}(\mathcal{E})$. ∎
+Now we prove Proposition 3.2.
 Proof.
-If there exists some $R\subseteq[m]$ such that $H_{0,R}(\mathcal{E})$ is not rejected, that is, $p_{R}>\alpha$.
-Then $\tilde{p}^{*}_{i}(\alpha)=\max\{p^{*}_{i},\Pi_{S\subseteq[m]}\mathbbm{1}_{p_{S}\leq\alpha}\}=\max\{p^{*}_{i},0\}=p^{*}_{i}$, so $\tilde{S}(\mathcal{E})=\widehat{S}(\mathcal{E})=\widehat{S}^{\text{ICP}}(\mathcal{E})$ by Lemma A.1.
+(Proposition 3.2) If there exists some $R\subseteq[m]$ such that $H_{0,R}(\mathcal{E})$ is not rejected, that is, $p_{R}>\alpha$.
+Then $\tilde{p}^{*}_{i}(\alpha)=\max\{p^{*}_{i},\Pi_{S\subseteq[m]}\mathbbm{1}_{p_{S}\leq\alpha}\}=\max\{p^{*}_{i},0\}=p^{*}_{i}$, so $\tilde{S}(\mathcal{E})=\widehat{S}(\mathcal{E})=\widehat{S}^{\text{ICP}}(\mathcal{E})$, where the second equality is by Lemma B.1.
 Now we consider the remaining case that $H_{0,S}(\mathcal{E})$ is rejected for all $S\subseteq[m]$, that is, $p_{S}\leq\alpha$ for all $S\subseteq[m]$.
@@
 ∎
-A.3 Proof of Proposition 3
-Proposition A.3.
-For $\alpha\in(0,1)$, we have $P\left(\widehat{S}(\mathcal{E})\subseteq S^{*}\right)\geq 1-\alpha.$
+B.3 Proof of Proposition 3.3
 Proof.
-When $H_{0,S^{*}}(\mathcal{E})$ is not rejected, for any $i\not\in S^{*}$, we have $p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}\geq p_{S^{*}}>\alpha$, so $i\not\in\widehat{S}(\mathcal{E})$.
+(Proposition 3.3) When $H_{0,S^{*}}(\mathcal{E})$ is not rejected, for any $i\not\in S^{*}$, we have $p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}\geq p_{S^{*}}>\alpha$, so $i\not\in\widehat{S}(\mathcal{E})$.
 Hence, we have $\widehat{S}(\mathcal{E})\subseteq S^{*}$.
@@
 $P\left(\widehat{S}(\mathcal{E})\subseteq S^{*}\right)\geq P\left(H_{0,S^{*}}(\mathcal{E})\text{ is not rejected}\right)\geq 1-\alpha.$
-A.4 Proof of Proposition 4
-Proposition A.4.
-For any $c\in(0,1)$, $P_{H^{*}_{0,S}}(p^{*}_{S}\leq c)\leq c$.
+B.4 Proof of Proposition 4.1
 Proof.
-When $H^{*}_{0,S}$ is true, we have $S^{*}\subseteq[m]\setminus S$ as $S\cap S^{*}=\emptyset$, so $p^{*}_{S}=\max_{I\subseteq[m]\setminus S}p_{I}\geq p_{S^{*}}$. Therefore, for any $c\in(0,1)$,
+(Proposition 4.1) When $H^{*}_{0,S}$ is true, we have $S^{*}\subseteq[m]\setminus S$ as $S\cap S^{*}=\emptyset$, so $p^{*}_{S}=\max_{I\subseteq[m]\setminus S}p_{I}\geq p_{S^{*}}$.
+Therefore, for any $c\in(0,1)$,
 $P_{H^{*}_{0,S}}(p^{*}_{S}\leq c)\leq P_{H^{*}_{0,S}}(p_{S^{*}}\leq c)\leq c.$
-A.5 Proof of Theorem 1
-Theorem A.1.
-Let $\alpha\in(0,1)$, then
-$P(|R\cap\mathcal{N}^{*}|\leq t_{\alpha}(R)\text{ for any }R\subseteq[m])\geq 1-\alpha.$
+B.5 Proof of Theorem 4.1
+This result directly follows from the fact that $t_{\alpha}(R)$ is equivalent to the simultaneous true discovery bound based on closed testing, as discussed in Section 4.2.
+We also give a self-contained proof below for completeness.
 Proof.
-For any $R\subseteq[m]$,
-$p^{*}_{R\cap\mathcal{N}^{*}}=\max_{I\subseteq[m]\setminus(R\cap\mathcal{N}^{*})}p_{I}\geq\max_{I\subseteq[m]\setminus\mathcal{N}^{*}}p_{I}=p^{*}_{\mathcal{N}^{*}}.$
-So if $p^{*}_{\mathcal{N}^{*}}>\alpha$, we have $p^{*}_{R\cap\mathcal{N}^{*}}>\alpha$, which implies that $|R\cap\mathcal{N}^{*}|\leq t_{\alpha}(R)$ by definition. Hence
-$P(|R\cap\mathcal{N}^{*}|\leq t_{\alpha}(R)\text{ for any }R\subseteq[m])\geq P(p^{*}_{\mathcal{N}^{*}}>\alpha).$
+(Theorem 4.1) If $p^{*}_{\mathcal{N}^{*}}>\alpha$, we have $|R\cap S^{*}|=|R\setminus\mathcal{N}^{*}|\geq t_{\alpha}(R)$ by definition.
+Hence
+$P(|R\cap S^{*}|\geq t_{\alpha}(R)\text{ for any }R\subseteq[m])\geq P(p^{*}_{\mathcal{N}^{*}}>\alpha).$
 In addition, we have
@@
 which completes the proof. ∎
-A.6 Proof of Proposition 5
-Proposition A.5.
-Let $S^{*}$ be the set of true causal predictors.
-Consider the desired case where only $H_{0,S^{*}}(\mathcal{E})$ is true, $p_{S^{*}}\sim U(0,1)$ and $p_{S}=0$ for any $S\neq S^{*}$.
-Then, for $p^{\prime}_{i}$ satisfying (15), there exists some $c^{*}\in(0,1)$ such that $P_{H^{*}_{0,i}}(p^{\prime}_{i}\leq c^{*})>c^{*}$.
+B.6 Proof of Theorem 4.2
 Proof.
-For any $i$ such that $H^{*}_{0,i}$ is true, we have $p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}=p_{S^{*}}$.
-Since $P_{H^{*}_{0,i}}(p^{\prime}_{i}<p^{*}_{i})>0$, there must exist some rational number $c^{*}\in(0,1)$ such that
-$P_{H^{*}_{0,i}}(p^{\prime}_{i}\leq c^{*}<p^{*}_{i})>0.$ (16)
-Then,
-$P_{H^{*}_{0,i}}(p^{\prime}_{i}\leq c^{*})$
-$=$ $P_{H^{*}_{0,i}}(p^{\prime}_{i}\leq c^{*},p^{\prime}_{i}<p^{*}_{i})+P_{H^{*}_{0,i}}(p^{\prime}_{i}\leq c^{*},p^{\prime}_{i}=p^{*}_{i})$
-$=$ $P_{H^{*}_{0,i}}(p^{\prime}_{i}\leq c^{*},p^{\prime}_{i}<p^{*}_{i}\leq c^{*})+P_{H^{*}_{0,i}}(p^{\prime}_{i}\leq c^{*}<p^{*}_{i})+P_{H^{*}_{0,i}}(p^{\prime}_{i}\leq c^{*},p^{\prime}_{i}=p^{*}_{i})$
-$>$ $P_{H^{*}_{0,i}}(p^{*}_{i}\leq c^{*},p^{\prime}_{i}<p^{*}_{i})+P_{H^{*}_{0,i}}(p^{*}_{i}\leq c^{*},p^{\prime}_{i}=p^{*}_{i})$
-$=$ $P_{H^{*}_{0,i}}(p^{*}_{i}\leq c^{*})$
-$=$ $P_{H^{*}_{0,i}}(p_{S^{*}}\leq c^{*})$
-$=$ $c^{*},$
-which completes the proof. ∎
+(Theorem 4.2) Let
+$u_{\alpha}(R)=|R|-t_{\alpha}(R)=\max\{|I|:I\subseteq R\text{ and}p^{*}_{I}>\alpha\}.$
+and
+$\tilde{u}_{\alpha}(R)=|R|-\tilde{t}_{\alpha}(R)=\underset{S\subseteq[m]\text{ with }p_{S}>\alpha}{\max}|R\setminus S|$
+Then, to show that $t_{\alpha}(R)=\tilde{t}_{\alpha}(R)$, it suffices to show that $u_{\alpha}(R)=\tilde{u}_{\alpha}(R)$.
+Let
+$A(R)=\{I\subseteq R:\max_{U\subseteq[m]\setminus I}p_{U}>\alpha\}\quad\text{and}\quad B=\{S\subseteq[m]:p_{S}>\alpha\},$
+then we have
+$u_{\alpha}(R)=\max\{|I|:I\subseteq R\text{ and }p^{*}_{I}>\alpha\}=\max\{|I|:I\in A(R)\}$
+and
+$\tilde{u}_{\alpha}(R)=\max\{|R\setminus S|:S\subseteq[m]\text{ and }p_{S}>\alpha\}=\max\{|R\setminus S|:S\in B\}.$
+Let $I_{1}\in A(R)$ and $S_{2}\in B$ be the sets for which $u_{\alpha}(R)$ and $\tilde{u}_{\alpha}(R)$ take their maximal values, respectively. That is,
+$u_{\alpha}(R)=|I_{1}|\quad\text{and}\quad\tilde{u}_{\alpha}(R)=|R\setminus S_{2}|=:|I_{2}|.$
+First, note that $I_{2}=R\setminus S_{2}\subseteq R$ implies that $S_{2}\subseteq[m]\setminus I_{2}$, so $\underset{U\subseteq[m]\setminus I_{2}}{\max}p_{U}\geq P_{S_{2}}>\alpha$, where the second inequality holds as $S_{2}\in B$.
+Hence, we have $I_{2}\in A(R)$, which implies that $|I_{2}|\leq|I_{1}|$ by the definition of $I_{1}$.
+To show that $u_{\alpha}(R)=\tilde{u}_{\alpha}(R)$, it is sufficient to prove that $|I_{1}|\leq|I_{2}|$.
+Let $S_{1}=\operatorname*{argmax}_{U\subseteq[m]\setminus I_{1}}p_{U}$, so $S_{1}\subseteq[m]\setminus I_{1}$.
+Together with the fact that $I_{1}\subseteq R$, we have $I_{1}\subseteq R\setminus S_{1}$, so $|I_{1}|\leq|R\setminus S_{1}|$.
+In addition, since $I_{1}\in A(R)$, we have $p_{S_{1}}>\alpha$.
+That is, $S_{1}\in B$.
+Hence, $|R\setminus S_{1}|\leq|R\setminus S_{2}|$ by the definition of $S_{2}$.
+Finally, we have $|I_{1}|\leq|R\setminus S_{1}|\leq|R\setminus S_{2}|=|I_{2}|$.
+∎